# SphereDrag: Spherical Geometry-Aware Panoramic Image Editing

### v1

@@ 4.2 Performance Comparison
Figure 7: Influence of different hyperparameters $t$ across various FOV settings (30, 60, and 90). The top row (a-c) shows the relationship between $\lambda$ and the IF value, while the bottom row (d-f) shows $\lambda$ versus FID and sFID for the corresponding FOV.
4.3 Ablation Studies
In this section, we analyze the contribution of each module to overall performance on images with FOV angles of 30, 60, and 90. The evaluation is conducted using the IF metric.
The results demonstrate that AR leads to consistent improvements over the baseline in all tested FOV settings. Adding the GCTA yields more improvements. When combining AR, GCTA, and SSRT, further improvements are achieved in all tested FOV settings. Also, we show some visualization for AR. As shown in fig. 6, AR substantially improves semantic continuity and geometric structure in output panoramic image.
Figure 8: PCA visualization results. At smaller timesteps (e.g., $t=1,10$), the latent variable contains contains redundant details, while at larger timesteps (e.g., $t=50$), the spatial information is overly diminished. At $t=35$, a good balance is achieved with fewer noise and richer spatial information, effectively preserving spatial coherence.
4.4 Discussion
Hyperparameters $t$ As shown in fig. 7, based on the evaluation in three FOVs (30, 60 and 90), $t=35$ demonstrates consistently strong performance in IF, FID and sFID, with noticeable improvements in narrower FOVs (30 and 60) and competitive results in wider FOV (90), indicating robust generalization. In comparison, configurations such as $t=30$, $t=40$, and $t=45$ display more variability and certain limitations in different settings. Furthermore, the PCA visualization in fig. 8 reveals that $t=35$ achieves a favorable balance between noise suppression and spatial information retention, offering latent representations that effectively preserve spatial coherence. Overall, $t=35$ emerges as a reasonable default choice, offering balanced and reliable performance across a wide range of FOV conditions.
@@ 4.4 Discussion
Figure 9: Influence of different hyperparameters $\lambda$ across various FOV settings (30, 60, and 90). The top row (a-c) shows the relationship between $\lambda$ and the IF value, while the bottom row (d-f) shows $\lambda$ versus FID and sFID for the corresponding FOV.
Hyperparameter $\lambda$ controls the proportion of the second term in the motion supervision loss. Specifically, we evaluate $\lambda$ values of 0, 0.05, 0.1, 0.15, and 0.2. When $\lambda=0$, the second term of the motion supervision loss is removed, leading to a considerable decrease in IF, FID and sFID, as shown in fig. 9, which highlights the importance of this loss term in improving the model’s performance. When $\lambda>0$, the performance generally improves, and is not sensitive to $\lambda$. The performance fluctuations remain small, demonstrating the robustness of our method with respect to $\lambda$. By comparing the performance across different FOV settings, we set $\lambda=0.1$ finally for its best and most stable performance across all conditions.
5 Conclusion
Appendix A Detail for $\mathcal{T}_{\text{align}}(\cdot)$ in Adaptive Reprojection (AR)
$\theta=\frac{\pi}{2}-\frac{j}{H}\cdot\pi,\quad\phi=\frac{i}{W}\cdot 2\pi-\pi.$ (14)
$R_{\phi}(\Delta\phi)=\begin{bmatrix}\cos(\Delta\phi)&-\sin(\Delta\phi)&0\\ \sin(\Delta\phi)&\cos(\Delta\phi)&0\\ 0&0&1\end{bmatrix},$ (15)
$R_{\theta}(\Delta\theta)=\begin{bmatrix}1&0&0\\ 0&\cos(\Delta\theta)&-\sin(\Delta\theta)\\ 0&\sin(\Delta\theta)&\cos(\Delta\theta)\end{bmatrix},$ (16)
$R(\Delta\theta,\Delta\phi)=R_{\theta}(\Delta\theta)\cdot R_{\phi}(\Delta\phi),$ (17)
Appendix B Detail for Great Circle Direction Calculation
$\theta=\frac{\pi}{2}-\frac{j}{H}\cdot\pi,\quad\phi=\frac{i}{W}\cdot 2\pi-\pi,$ (18)
$x$ $=\cos(\theta)\cos(\phi),$ (19)
$y$ $=\cos(\theta)\sin(\phi),$
$z$ $=\sin(\theta).$
$\mathbf{n}=\frac{\mathbf{P}_{\text{han}}\times\mathbf{P}_{\text{tar}}}{\|\mathbf{P}_{\text{han}}\times\mathbf{P}_{\text{tar}}\|}.$ (20)
$\mathbf{P}^{\prime}=\frac{\mathbf{P}_{\text{han}}^{k}-(\mathbf{P}_{\text{han}}^{k}\cdot\mathbf{n})\mathbf{n}}{\|\mathbf{P}_{\text{han}}^{k}-(\mathbf{P}_{\text{han}}^{k}\cdot\mathbf{n})\mathbf{n}\|}.$ (21)
$\mathbf{v}_{\text{move}}=\frac{\mathbf{P}_{\text{tar}}-\mathbf{P}^{\prime}}{\|\mathbf{P}_{\text{tar}}-\mathbf{P}^{\prime}\|}.$ (22)
[…]
$\hat{\mathbf{e}}_{\theta}$ $=\frac{\partial\mathbf{P}}{\partial\theta}=\left[-\sin(\theta)\cos(\phi),\ -\sin(\theta)\sin(\phi),\ \cos(\theta)\right]^{T}.$ (24)
$\begin{bmatrix}v_{\phi}\\ v_{\theta}\par\end{bmatrix}=\begin{bmatrix}\mathbf{v}_{\text{move}}\cdot\hat{\mathbf{e}}_{\phi}\\ \mathbf{v}_{\text{move}}\cdot\hat{\mathbf{e}}_{\theta}\par\end{bmatrix}.$ (25)
$\vec{d}=\frac{[\frac{v_{\phi}}{\cos(\theta)},v_{\theta}]}{\left\|\left[\frac{v_{\phi}}{\cos(\theta)},v_{\theta}\right]\right\|}.$ (26)

### v2

@@ E Evaluation Metrics
Figure S-3: Influence of different hyperparameters $t$ across various FOV settings (30, 60, and 90). The top row (a-c) shows the relationship between $\lambda$ and the IF value, while the bottom row (d-f) shows $\lambda$ versus FID and sFID for the corresponding FOV.
@@
Figure S-4: Influence of different hyperparameters $\lambda$ across various FOV settings (30, 60, and 90). The top row (a-c) shows the relationship between $\lambda$ and the IF value, while the bottom row (d-f) shows $\lambda$ versus FID and sFID for the corresponding FOV.
F Discussion
Hyperparameters $t$ As shown in fig. S-3, based on the evaluation in three FOVs (30, 60 and 90), $t=35$ demonstrates consistently strong performance in IF, FID and sFID, with noticeable improvements in narrower FOVs (30 and 60) and competitive results in wider FOV (90), indicating robust generalization. In comparison, configurations such as $t=30$, $t=40$, and $t=45$ display more variability and certain limitations in different settings. Furthermore, the PCA visualization in fig. S-2 reveals that $t=35$ achieves a favorable balance between noise suppression and spatial information retention, offering latent representations that effectively preserve spatial coherence. Overall, $t=35$ emerges as a reasonable default choice, offering balanced and reliable performance across a wide range of FOV conditions.
Hyperparameter $\lambda$ controls the proportion of the second term in the motion supervision loss. Specifically, we evaluate $\lambda$ values of 0, 0.05, 0.1, 0.15, and 0.2. When $\lambda=0$, the second term of the motion supervision loss is removed, leading to a considerable decrease in IF, FID and sFID, as shown in fig. S-4, which highlights the importance of this loss term in improving the model’s performance. When $\lambda>0$, the performance generally improves, and is not sensitive to $\lambda$. The performance fluctuations remain small, demonstrating the robustness of our method with respect to $\lambda$.
By comparing the performance across different FOV settings, we set $\lambda=0.1$ finally for its best and most stable performance across all conditions.
G Effectiveness of Adaptive Reprojection
htb] Refer to caption
Figure S-5: Ablation study on adaptive reprojection (AR). The top row shows the result with AR enabled, while the bottom row shows the result without it. The red boxes highlight the seam regions. With AR, the transitions across the seams are smooth and visually coherent. In contrast, without AR, noticeable artifacts and discontinuities appear. These results demonstrate the effectiveness of AR in enhancing geometric alignment and visual consistency in panoramic image editing.